# Monotonicity of the ultrafilter number function

### v3

@@ 7. Monotonicity at a singular cardinal and its successor
We may assume $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$. Fix an increasing sequence $\langle{\kappa_{i}\mid i<\mathord{\mathrm{cf}}(\kappa)}\rangle$ of regular cardinals with limit $\kappa$. By Proposition 7.15, there is $\lambda<\kappa$ such that $\mathfrak{u}(\mu)\leq\mathfrak{u}(\kappa^{+})$ for every regular $\mu$ with $\lambda<\mu<\kappa$. Hence we may assume that for every $i<\mathord{\mathrm{cf}}(\kappa)$, there is a uniform ultrafilter $\mathcal{U}_{i}$ over $\kappa_{i}$ with $\chi(\mathcal{U}_{i})\leq\mathfrak{u}(\kappa^{+})$. Fix also a uniform ultrafilter $\mathcal{V}$ over $\mathord{\mathrm{cf}}(\kappa)$. Now define a filter $\mathcal{D}$ over $\kappa$ as: $X\in\mathcal{D}\iff\{i<\mathord{\mathrm{cf}}(\kappa)\mid X\cap\kappa_{i}\in\mathcal{U}_{i}\}\in\mathcal{V}$. It is easy to show that $\mathcal{D}$ is a uniform ultrafilter over $\kappa$, and since $\chi(\mathcal{U}_{i})\leq\mathfrak{u}(\kappa^{+})$, we have $\chi(\mathcal{D})\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}\cdot 2^{\mathord{\mathrm{cf}}(\kappa)}$. Thus $\mathfrak{u}(\kappa)\leq\chi(\mathcal{D})\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}$. ∎
@@
We prove that $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mathfrak{u}(\kappa^{+\omega_{1}+1})$ in the generic extension. Suppose to the contrary that $\mathfrak{u}(\kappa^{+\omega_{1}+1})<\mathfrak{u}(\kappa^{+\omega_{1}})$. We note that $\mathfrak{u}(\kappa^{+\omega_{1}+1})=\mu$ by Lemma 4.1. By Lemma 7.17, we have $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mathfrak{u}(\kappa^{+\omega_{1}+1})^{\omega_{1}}$. However, since $\mathfrak{u}(\kappa^{+\omega_{1}+1})=\mu$, $\mathord{\mathrm{cf}}(\mu)=\omega_{3}$, and $\mathrm{Add}(\omega_{2},\mu)$ is $\omega_{2}$-closed, we have $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mu^{\omega_{1}}=\mu=$ $\mathfrak{u}(\kappa^{+\omega_{1}+1})$, this is a contradiction. ∎
@@ 8. Consistency strength of the failure of monotonicity
In this section, we study the consistency strength of the failure of monotonicity of the ultrafilter number function. We will show that the failure has a large cardinal strength.
@@
Fact 8.1 (Donder [4]).
Suppose there is no inner model of a measurable cardinal. Let $\kappa$ be an uncountable cardinal.
@@
Suppose there is a cardinal $\kappa$ such that $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter for some $\lambda<\kappa$. Then there is an inner model of a measurable cardinal.
@@
Let $\mathcal{U}$ be a $\lambda$-indecomposable uniform ultrafilter over $\kappa$. Then $\mathcal{U}$ is not $(\omega,\lambda)$-regular by Lemma 2.7, and the assertion follows from Fact 8.1. ∎
@@
Suppose there are cardinals $\lambda<\kappa$ such that $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$. Then there is an inner model of a measurable cardinal.
@@
By the assumption, $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter. Then there is an inner model of a measurable cardinal by Proposition 8.2. ∎
@@
Relative to a certain large cardinal assumption, it is consistent that there is a weakly inaccessible cardinal $\kappa$ such that $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$ for every uncountable cardinal $\lambda<\kappa$.
@@
Suppose there is a measurable cardinal $\kappa$ and a normal measure $U$ over $\kappa$ such that $V=L[U]$, and $V_{\kappa}$ is a model of $\mathsf{ZFC}+$“there is no inner model of a measurable cardinal”. By Fact 8.1, for every uncountable cardinal $\lambda<\kappa$, $\lambda$ has no $\omega_{1}$-indecomposable uniform ultrafilter.
Fix a large strong limit singular cardinal $\mu>\kappa$ with cofinality $\omega_{1}$, and take a generic extension $V[G]$ via poset $\mathrm{Add}(\omega,\mu)$. In $V[G]$, we have $\mathfrak{u}(\kappa)\leq\mu<\mathfrak{u}(\omega_{1})$. We see that, in $V[G]$, every uncountable cardinal $\lambda<\kappa$ does not have an $\omega_{1}$-indecomposable uniform ultrafilter. Then $\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\lambda)$, hence $\mathfrak{u}(\kappa)<\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\lambda)$ as required. Suppose to the contrary that $\lambda$ has an $\omega_{1}$-indecomposable uniform ultrafilter. Since $V[G]$ is a c.c.c. forcing extension of $V$, by the argument used in the proof of Theorem 6.9, we can take an $\omega_{1}$-indecomposable uniform ultrafilter over $\lambda$ in $V$. This contradicts to the choice of $V$. ∎
@@
Clearly $\square_{\kappa}\Rightarrow\square(\kappa^{+})$. It is known that the failure of the square principle at a singular cardinal, or at successive cardinals have a very strong large cardinal strength, but the exact strength is unknown. We just present the following Schimmerling’s result.
Fact 8.8 (Schimmerling [18]).
@@
Then there is an inner model of a proper class of strong cardinals.
@@
For this sake, we need the next definition and Kanamori’s another result.
@@
    A function $f:\kappa\to\kappa+1$ is a least function of $\mathcal{U}$ if $\{\alpha<\kappa\mid\beta<f(\alpha)\}\in\mathcal{U}$ for every $\beta<\kappa$, and for every $g:\kappa\to\kappa$, if $\{\alpha\mid g(\alpha)<f(\alpha)\}\in\mathcal{U}$, then there is $\beta<\kappa$ such that $\{\alpha\mid g(\alpha)\leq\beta\}\in\mathcal{U}$.
@@
To show that $\square(\kappa)$ fails, suppose to the contrary that there is a $\square(\kappa)$-sequence $\langle{c_{\alpha}\mid\alpha<\kappa}\rangle$. By Fact 8.11, $\mathcal{U}$ has a least function. Hence we may assume that $\mathcal{U}$ is weakly normal.
@@
(1). By the assumption, we can find a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$. If $\lambda$ is regular, then $\square(\kappa)$ fails by Proposition 8.9. If $\lambda$ is singular, then $\mathcal{U}$ is either $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, or $\lambda^{+}$-indecomposable by Proposition 7.3. In either case, we have that $\square(\kappa)$ fails by Proposition 8.9 again. ∎
The failure of the square principle is reflecting.
@@
By the assumption, $\kappa^{+}$ has a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$. We know that $\mathcal{U}$ is $\kappa$-decomposable. Now we show that there is a regular cardinal $\mu<\kappa$ such that $\mathcal{U}$ is $\mu$-indecomposable. Since $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$, there is a $\lambda$-indecomposable uniform ultrafilter. If $\lambda$ is regular, let $\mu=\lambda$. Suppose $\lambda$ is singular. By Proposition 7.3, $\mathcal{U}$ is either $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, or else $\lambda^{+}$-indecomposable. If it is $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, then let $\mu=\mathord{\mathrm{cf}}(\lambda)$. If $\lambda^{+}$-indecomposable, we have $\lambda^{+}<\kappa$ and put $\mu=\lambda^{+}$.
We have that $\square(\kappa^{+})$ fails by Proposition 8.9. Because $\mathcal{U}$ is $\kappa$-decomposable, we can find a $\mu$-indecomposable uniform ultrafilter over $\kappa$. By applying Proposition 8.9 again, $\square(\kappa)$ fails. ∎
@@
By the assumption, we can take a $\kappa$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa^{+}$, which is in fact $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Proposition 7.3. Now suppose to contrary that the set $\{\lambda<\kappa\mid\lambda$ is regular and $\square(\lambda)$ holds }}\}} is unbounded in $\kappa$. By Proposition 7.15, we can take a regular $\lambda<\kappa$ such that $\lambda>\mathord{\mathrm{cf}}(\kappa)$, $\mathcal{U}$ is $\lambda$-decomposable, and $\square(\lambda)$ holds. Then we can take a $\mathord{\mathrm{cf}}(\kappa)$-indecomposable uniform ultrafilter $\mathcal{V}$ over $\lambda$. Hence $\square(\lambda)$ fails by Proposition 8.9, this is a contradiction. ∎
By Fact 8.8 and these propositions, we obtain the following lower bound of the consistency strength.
@@
Let $\kappa$ be a cardinal. If $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$ holds for some cardinal $\lambda\leq\kappa$, then there is an inner model of a proper class of strong cardinals.
@@
    What is the exact consistency strength of the statement that “$\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$ for some $\lambda\leq\kappa$”?
@@
    Is there anther cardinal invariant function $\mathfrak{k}(\kappa)$ on the cardinals such that the failure of monotonicity of $\mathfrak{k}$ has a large cardinal strength?
Acknowledgements
The author thanks Assaf Rinot for many useful comments.

### v4

@@ 7. Monotonicity at a singular cardinal and its successor
It is trivial if $\kappa$ is regular, so suppose $\kappa$ is singular. We may assume $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\kappa)$. Fix an increasing sequence $\langle{\kappa_{i}\mid i<\mathord{\mathrm{cf}}(\kappa)}\rangle$ of regular cardinals with limit $\kappa$. By Theorem 7.15, there is $\lambda<\kappa$ such that $\mathfrak{u}(\mu)\leq\mathfrak{u}(\kappa^{+})$ for every regular $\mu$ with $\lambda<\mu<\kappa$. Hence we may assume that for every $i<\mathord{\mathrm{cf}}(\kappa)$, there is a uniform ultrafilter $\mathcal{U}_{i}$ over $\kappa_{i}$ with $\chi(\mathcal{U}_{i})\leq\mathfrak{u}(\kappa^{+})$. Fix also a uniform ultrafilter $\mathcal{V}$ over $\mathord{\mathrm{cf}}(\kappa)$. Now define a filter $\mathcal{D}$ over $\kappa$ as: $X\in\mathcal{D}\iff\{i<\mathord{\mathrm{cf}}(\kappa)\mid X\cap\kappa_{i}\in\mathcal{U}_{i}\}\in\mathcal{V}$. It is easy to show that $\mathcal{D}$ is a uniform ultrafilter over $\kappa$, and since $\chi(\mathcal{U}_{i})\leq\mathfrak{u}(\kappa^{+})$, we have $\chi(\mathcal{D})\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}\cdot 2^{\mathord{\mathrm{cf}}(\kappa)}$. Thus $\mathfrak{u}(\kappa)\leq\chi(\mathcal{D})\leq\mathfrak{u}(\kappa^{+})^{\mathord{\mathrm{cf}}(\kappa)}$. ∎
@@
We prove that $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mathfrak{u}(\kappa^{+\omega_{1}+1})$ in the generic extension. We know that $\mathfrak{u}(\kappa^{+\omega_{1}+1})\geq\mu$ by Lemma 4.1. By Lemma 7.16, we have $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mathfrak{u}(\kappa^{+\omega_{1}+1})^{\omega_{1}}$. However, since $\mathfrak{u}(\kappa^{+\omega_{1}+1})=\mu$, $\mathord{\mathrm{cf}}(\mu)=\omega_{3}$, and $\mathrm{Add}(\omega_{2},\mu)$ is $\omega_{2}$-closed, we have $\mathfrak{u}(\kappa^{+\omega_{1}})\leq\mu^{\omega_{1}}=\mu=$ $\mathfrak{u}(\kappa^{+\omega_{1}+1})$. ∎
@@ 8. Consistency strength of the failure of monotonicity
In this section, we study the consistency strength of the failure of monotonicity of the ultrafilter number function. We will show that the failure has large cardinal strength.
@@
Theorem 8.1 (Theorem 4.1 and 4.5 in Donder [6]).
Suppose there is no inner model with a measurable cardinal. Let $\kappa$ be an uncountable cardinal.
@@
Suppose there is a cardinal $\kappa$ such that $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter for some $\lambda<\kappa$. Then there is an inner model with a measurable cardinal.
@@
Let $\mathcal{U}$ be a $\lambda$-indecomposable uniform ultrafilter over $\kappa$. Then $\mathcal{U}$ is not $(\omega,\lambda)$-regular by Lemma 2.7. If $\kappa$ is regular, by (2) of Theorem 8.1, there is an inner model with a measurable cardinal. Suppose $\kappa$ is singular. Since $\lambda<\kappa$ and $U$ is not $(\omega,\lambda)$-regular, it is easy to check that $U$ is not $(\omega,\kappa)$-regular. thus there is an inner model with a measurable cardinal by (1) of Theorem 8.1. ∎
@@
Suppose there are cardinals $\lambda<\kappa$ such that $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$. Then there is an inner model with a measurable cardinal.
@@
By the assumption Corollary 5.2, $\kappa$ has a $\lambda$-indecomposable uniform ultrafilter. Then there is an inner model with a measurable cardinal by Proposition 8.2. ∎
@@
Relative to the existence of a measurable cardinal, it is consistent that there is a weakly inaccessible cardinal $\kappa$ such that $\mathfrak{u}(\kappa)<\mathfrak{u}(\lambda)$ for every uncountable cardinal $\lambda<\kappa$.
@@
Suppose there is a measurable cardinal $\kappa$ and a normal measure $U$ over $\kappa$ such that $V=L[U]$, and $V_{\kappa}$ is a model with $\mathsf{ZFC}+$“there is no inner model with a measurable cardinal”. By Theorem 8.1 applied in $V_{\kappa}$, for every uncountable cardinal $\lambda<\kappa$, $\lambda$ has no $\omega_{1}$-indecomposable uniform ultrafilter.
Fix a large strong limit singular cardinal $\mu>\kappa$ with cofinality $\omega_{1}$, and take a generic extension $V[G]$ via poset $\mathrm{Add}(\omega,\mu)$. In $V[G]$, we have $\mathfrak{u}(\kappa)\leq\mu<\mathfrak{u}(\omega_{1})$. It is sufficient to show that $\mathfrak{u}(\omega_{1})\leq\mathfrak{u}(\lambda)$. Suppose to the contrary that $\mathfrak{u}(\omega_{1})>\mathfrak{u}(\lambda)$. Then $\lambda$ has an $\omega_{1}$-indecomposable uniform ultrafilter by Corollary 5.2. Since $V[G]$ is a c.c.c. forcing extension of $V$, by the argument used in the proof of Theorem 6.11, we can take an $\omega_{1}$-indecomposable uniform ultrafilter over $\lambda$ in $V$. This contradicts to the choice of $V$. ∎
@@
Clearly $\square_{\kappa}\Rightarrow\square(\kappa^{+})$. It is known that the failure of the square principle at a singular cardinal, or at successive cardinals have very strong large cardinal strength, but the exact strength is unknown. For example, the following theorem gives such a lower bound.
Theorem 8.8 (Schimmerling [22]).
@@
Then there is an inner model with a proper class of strong cardinals.
@@
For this sake, we need the next definition and yet another result of Kanamori.
@@
    A function $f:\kappa\to\kappa$ is a least function of $\mathcal{U}$ if $\{\alpha<\kappa\mid\beta<f(\alpha)\}\in\mathcal{U}$ for every $\beta<\kappa$, and for every $g:\kappa\to\kappa$, if $\{\alpha\mid g(\alpha)<f(\alpha)\}\in\mathcal{U}$, then there is $\beta<\kappa$ such that $\{\alpha\mid g(\alpha)\leq\beta\}\in\mathcal{U}$.
@@
To show that $\square(\kappa)$ fails, suppose to the contrary that there is a $\square(\kappa)$-sequence $\langle{c_{\alpha}\mid\alpha<\kappa}\rangle$. By Theorem 8.11, $\mathcal{U}$ has a least function. Hence we may assume that $\mathcal{U}$ is weakly normal.
@@
(1). By the assumption, we can find a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$. If $\lambda$ is regular, then $\square(\kappa)$ fails by Proposition 8.9. If $\lambda$ is singular, then $\mathcal{U}$ is either $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, or $\lambda^{+}$-indecomposable by Theorem 7.3. In either case, we have that $\square(\kappa)$ fails by Proposition 8.9 again. ∎
@@
First note that in fact $\lambda\neq\kappa$ by Proposition 5.4. By the assumption, $\kappa^{+}$ has a $\lambda$-indecomposable uniform ultrafilter $\mathcal{U}$ by Corollary 5.2. Now we show that there is a regular cardinal $\mu<\kappa$ such that $\mathcal{U}$ is $\mu$-indecomposable. If $\lambda$ is regular, just let $\mu=\lambda$. Suppose $\lambda$ is singular. By Theorem 7.3, $\mathcal{U}$ is either $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, or else $\lambda^{+}$-indecomposable. If it is $\mathord{\mathrm{cf}}(\lambda)$-indecomposable, then let $\mu=\mathord{\mathrm{cf}}(\lambda)$. If it is $\lambda^{+}$-indecomposable, we have $\lambda^{+}<\kappa$ by Lemma 2.5. Put $\mu=\lambda^{+}$ in this case.
We have that $\square(\kappa^{+})$ fails by Proposition 8.9. $\mathcal{U}$ is $\kappa$-decomposable by Lemma 2.5, so we can find a $\mu$-indecomposable uniform ultrafilter over $\kappa$ by Lemma 2.8. By applying Proposition 8.9 again, $\square(\kappa)$ fails. ∎
@@
By the assumption, we can take a $\kappa$-indecomposable uniform ultrafilter $\mathcal{U}$ over $\kappa^{+}$, which is in fact $\mathord{\mathrm{cf}}(\kappa)$-indecomposable by Theorem 7.3. Now suppose to contrary that the set $\{\lambda<\kappa\mid\lambda$ is regular and $\square(\lambda)$ holds }\} is unbounded in $\kappa$. By Proposition 7.12, we can take a regular $\lambda<\kappa$ such that $\lambda>\mathord{\mathrm{cf}}(\kappa)$, $\mathcal{U}$ is $\lambda$-decomposable, and $\square(\lambda)$ holds. Then we can take a $\mathord{\mathrm{cf}}(\kappa)$-indecomposable uniform ultrafilter $\mathcal{V}$ over $\lambda$ by Lemma 2.8. Hence $\square(\lambda)$ fails by Proposition 8.9, this is a contradiction. ∎
By Theorem 8.8 and these propositions, we obtain the following lower bound of the consistency strength.
@@
Let $\kappa$ be a cardinal. If $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$ holds for some cardinal $\lambda\leq\kappa$, then there is an inner model with a proper class of strong cardinals.
@@
    What is the exact consistency strength of the statement $\mathfrak{u}(\kappa^{+})<\mathfrak{u}(\lambda)$ for some $\lambda\leq\kappa$”?
@@
    Is there anther cardinal invariant function $\mathfrak{k}(\kappa)$ on the cardinals such that the failure of monotonicity of $\mathfrak{k}$ has large cardinal strength?
Acknowledgments
The author deeply grateful to the referee for many insightful suggestions and comments. The author also would like to thank Assaf Rinot for many useful comments.